# Distributionally Robust Markov Games with Average Reward

### v2

@@ 4.2 Existence of a Nash Equilibrium
For any distributionally robust MDP, its optimal robust policy set $\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$ is convex.
As $\text{BR}_{i}(\pi_{-i})$ is the set of optimal robust policies of the (induced) distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, this result hence directly implies the convexity of $\text{BR}_{i}(\pi_{-i})$. We highlight that this proof is not a direct extension from its counterpart under the discounted reward as we need to tackle the non-uniqueness of solutions, whereas the discounted Bellman equation admits a unique solution.
@@
For convergent policy sequences $\{\pi^{n}_{-i}\}\to\pi_{-i}$ and $\{\mu^{n}_{i}\}\to\mu_{i}$, if $\mu^{n}_{i}$ is an optimal robust policy of the induced distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi^{n}_{-i}},r^{\pi^{n}_{-i}})$ for agent $i\in\mathcal{N}$, then $\mu_{i}$ is also optimal for $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$.
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, thus resulting in a distributionally robust MDP sequence with different dynamics. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
@@
The proofs of these results are based on the correspondence between the optimal policies and the solutions to the Bellman equations, heavily relying on Theorem˜4.1. In the next section when we study a weaker setting, such correspondence does not hold anymore and the proof derived cannot be applied there.
@@
(Existence of a Nash Equilibrium) For an average-reward distributionally robust Markov game satisfying Assumption 1, a stationary Nash equilibrium exists.
@@ 5 NE of Weakly Communicating DR-MGs
In this section, we further extend our studies to a more general setting of weakly communicating DR-MGs.
@@
$\mathcal{P}$ is convex and compact. For any transition kernel $\mathsf{P}\in\mathcal{P}$, and for any player $i$, the state space can be partitioned into two distinct sets $\mathcal{S}=\mathcal{S}_{0}^{i}\cup\mathcal{S}_{1}^{i}$, such that (1). Any state in $\mathcal{S}_{0}^{i}$ is transient under every stationary joint policy; And (2). For any two states $s,s^{\prime}\in\mathcal{S}_{1}^{i}$, and any policy $\pi_{-i}$, there exists some policy $\pi_{i}$ and $N$ such that $\mathbb{P}(S_{N}=s^{\prime}|S_{0}=s,(\pi_{i},\pi_{-i}),\mathsf{P})>0$.
@@
This assumption is adapted from the non-robust settings (wikecek2015stationary; bertsekas2011dynamic). The weakly communicating assumption is strictly weaker than ˜1, as it only requires accessibility under some policy, instead of every policy. Moreover, under this assumption, only the optimal robust average reward $g^{\star}_{\mathcal{P}}(s)$ is a constant over initial states, whereas $g^{\pi}_{\mathcal{P}}(s)$ is not.
The solvability of the optimal robust Bellman equation is derived recently in (wang2025bellman), as follows.
Lemma 5.1.
(wang2025bellman) Under ˜2, (8) has a solution $(g^{\star},u^{\star})$, and $g^{\star}=g^{\star}_{\mathcal{P}}(\rho)$.
The results are strictly weaker than the ones we derived for irreducible DR-MGs. Firstly, Theorem˜4.1 ensures the solvability of the robust Bellman equation with any policy, whereas Lemma˜5.1 only consider the optimal equation. Hence, a correspondence between the optimal robust policy and solutions of (8) cannot be developed as in irreducible settings, and the previous proofs cannot be applied to show the convexity and continuity of the best-response mapping.
To address this issue, we instead study another mapping specified through the optimal robust Bellman equation: for $s\in\mathcal{S}$, $\nu\in\Delta(\mathcal{A}_{i})$, and a joint policy $\pi_{-i}$, we define
$\Psi_{s}(\nu)\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\nu(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+\sigma_{(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}}(u^{\star})\big),$
and set a map as the solution set of the Bellman equation:
$\mathcal{G}_{i}(\pi_{-i})\ \triangleq\ \bigl\{\nu:\Psi_{s}(\nu(\cdot|s))=\max_{\mu\in\Delta(\mathcal{A}_{i})}\Psi_{s}(\mu),\forall s\in\mathcal{S}\bigr\}.$
Based on the solvability result in Lemma˜5.1, we can show that $\mathcal{G}_{i}$ is nonempty. Different from irreducible cases, due to the existence of transient states, it is not guaranteed that an optimal policy can be obtained from a solution to the Bellman equations, hence $\mathcal{G}_{i}(\pi_{-i})$ may not contain all best-response policies. However, we can show that $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$, which is sufficient to derive the existence of the NE. Further, if the map $\mathcal{G}(\pi)=\times_{i}\mathcal{G}_{i}(\pi_{-i})$ satisfies the conditions in Lemma˜4.2, it then admits a fixed point which is also a fixed point under the BR mapping, and hence implies the existence of NE. Our results are presented in the following theorem.
Theorem 5.1.
Under ˜2, the following statements hold. (1). $\mathcal{G}_{i}$ is nonempty, compact, and convex; (2). $\mathcal{G}_{i}$ is upper hemi-continuous; (3). $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$; (4). A DR-MG always has a stationary NE.
The proofs are deferred to Appendix˜D, which rely on the strong duality of the robust average reward shown in (wang2025bellman), and are fundamentally different from the ones for Section˜4.
@@ 6 Algorithmic Solutions
We then develop an algorithmic solution for average-reward DR-MGs. Notably, finding stationary NE is significantly challenging in general-sum games, even under non-robust settings (hu2003nash; li2007reinforcement; li2003learning; zinkevich2005cyclic), and existing convergence results generally require additional structural assumptions on the underlying games. In this section, we design a value iteration based algorithm in Algorithm 1, named robust Nash-Iteration, to find a NE. The convergence guarantee of our algorithm, similarly to previous ones for general-sum games, is only obtained under some additional assumptions. Our algorithm is designed for irreducible DR-MGs, as learning under weakly communicating settings is generally challenging and remains open (sahabandu2024rl).
@@
Initialize: $h_{i}=0,\forall i\in\mathcal{N}$
@@
return $\pi$
Our algorithm maintains an estimation, $h_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). This assumption is not merely a technical convenience; it is the conceptual linchpin that stabilizes the learning dynamics of the system. By placing constraints on which Nash Equilibrium the oracle returns, we ensure that the sequence of generated policies and value functions behaves predictably, guiding the algorithm steadily toward a solution instead of allowing it to wander chaotically or diverge.
@@
Under ˜1, and with a specific equilibrium selection oracle NE (see Assumption 3 in Appendix), the robust Nash-Iteration algorithm converges to a Nash Equilibrium.
This theorem provides the algorithmic contribution of our work, demonstrating that the proposed Robust Nash-Iteration algorithm is a principled method with guaranteed convergence. Given the known difficulties associated with multi-agent learning algorithms in general-sum games under the long-term performance criterion, our algorithm stands as a significant contribution. Taken together, our results provide a complete theoretical foundation for this problem. From establishing the existence of a robust equilibrium in Theorem 4.4 to now providing a concrete algorithm to find it, this work makes the class of average-reward DR-MGs amenable to computational analysis. It provides a blueprint for developing practical, reliable multi-agent systems that can find stable operating points even when faced with significant environmental uncertainty.

### v3

@@ 4.2 Existence of a Nash Equilibrium
For any distributionally robust MDP under ˜1, the optimal policy set $\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$ is convex.
As $\text{BR}_{i}(\pi_{-i})$ is the set of optimal robust policies of the (induced) distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, this result hence directly implies the convexity of $\text{BR}_{i}(\pi_{-i})$. We highlight that this proof is not a direct extension from its counterpart under the discounted reward as we need to tackle the non-contraction of the Bellman equation, whereas the discounted Bellman equation is a contraction and admits a unique solution. Also, compared to non-robust studies, we need to additionally tackle the non-linearity and uncertainty transition kernel introduced by the worst-case consideration. We also highlight that our proof heavily relies on the solvability and uniqueness (up to constant vectors) of (6), which enables us to fully characterize the optimal policies through the optimal robust Bellman equation.
@@
Under ˜1. For convergent policy sequences $\{\pi^{n}_{-i}\}\to\pi_{-i}$ and $\{\mu^{n}_{i}\}\to\mu_{i}$, if $\mu^{n}_{i}$ is an optimal robust policy of the induced distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi^{n}_{-i}},r^{\pi^{n}_{-i}})$ for agent $i\in\mathcal{N}$, then $\mu_{i}$ is also optimal for $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$.
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, resulting in a distributionally robust MDP sequence with different dynamics. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
@@
The proofs of these results rely on Theorem˜4.1, i.e., the correspondence between the optimal policies and the solutions to the Bellman equations, and the fact that the Bellman equation for any policy $\pi$ has a unique solution. In the next section where we study a weaker setting, such results do not hold anymore, and the proof derived cannot be applied there.
@@
(Existence of a Nash Equilibrium) For an average-reward DR-MG satisfying Assumption 1, a stationary robust Nash equilibrium exists.
@@ 5 NE of Weakly Communicating DR-MGs
In this section, we extend our existence result from irreducible DR-MGs to a more general class of weakly communicating DR-MGs. The weakly communicating assumption is known to be the weakest standard structural assumption under which efficient algorithms and sharp structural results are available for average-reward control; see, e.g., (wikecek2015stationary; bertsekas2011dynamic; wan2021learning).
@@
The uncertainty set $\mathcal{P}$ is convex and compact. Moreover, for any transition kernel $P\in\mathcal{P}$ and any player $i$, the state space can be decomposed as $\mathcal{S}=\mathcal{S}^{i}_{0}\cup\mathcal{S}^{i}_{1},$ with $\mathcal{S}^{i}_{0}\cap\mathcal{S}^{i}_{1}=\emptyset,$ such that: (1). Every state in $\mathcal{S}^{i}_{0}$ is transient under every stationary joint policy; (2). For any two states $s,s^{\prime}\in\mathcal{S}^{i}_{1}$ and any stationary policy $\pi_{-i}$ of the other players, there exist a stationary policy $\pi_{i}$ for player $i$ and an integer $N$ (possibly depending on $s,s^{\prime},\pi_{-i},P$) such that $\mathbb{P}\big(S_{N}=s^{\prime}\mid S_{0}=s,(\pi_{i},\pi_{-i}),P\big)>0.$
@@
As in the non-robust case, weakly communicating is strictly weaker than irreducibility: it only requires accessibility under some policy, rather than under all joint policies. Under Assumption 2, only the optimal robust average reward $g^{\star}_{\mathcal{P}}(s)$ is constant across initial states, whereas the robust average reward $g^{\pi}_{\mathcal{P}}(s)$ of a general stationary policy $\pi$ may depend on $s$; see, e.g., (puterman2014markov; atia2021steady) for the non-robust case and wang2025bellman for the robust extension.
Moreover, the solvability guarantees for weakly communicating robust Bellman equations are weaker. The solvability of the optimal robust Bellman equation under the weakly communicating assumption was recently established by wang2025bellman.
Lemma 5.1 (Optimal robust Bellman equation (wang2025bellman)).
Let an average-reward DR-MDP satisfy Assumption 2. Then the optimal robust Bellman equation
$V(s)\;=\;\max_{a\in\mathcal{A}}\Big\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\Big\},\qquad s\in\mathcal{S},$ (8)
admits a solution $(g^{\star},u^{\star})$, with $g^{\star}=g^{\star}_{\mathcal{P}}(\rho)$ equal to the optimal robust average reward for any initial distribution $\rho$. Moreover, $g^{\star}$ is the optimal min–max value $\sup_{\pi}\inf_{P\in\mathcal{P}}g^{\pi}_{P}$, and strong min–max duality holds.
Compared with the irreducible case (Theorem 4.1), these results are strictly weaker. Theorem 4.1 gives solvability of the robust Bellman equation for every stationary policy $\pi$, as well as a tight correspondence between optimal policies and solutions of the optimal Bellman equation. Under Assumption 2, Lemma 5.1 guarantees solvability only for the optimal equation (8); the equation associated with a fixed policy $\pi$ may fail to have a solution, and one can no longer characterize optimal policies solely via the Bellman equations. As a consequence, the proof strategy of irreducible DR-MGs cannot be directly reused.
To address this issue, we construct a proxy set-valued map, derived from solutions of the optimal robust Bellman equation, and show that this map is contained in the true best-response map. Fix a player $i$ and a stationary joint policy $\pi_{-i}$ of all other players, and consider the induced single-agent DR-MDP $M_{i}(\pi_{-i})=(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i})$. Applying Lemma 5.1 to $M_{i}(\pi_{-i})$ yields a solution $(g^{\star}_{\pi_{-i}},u^{\star}_{\pi_{-i}})$ of the optimal Bellman equation for this induced DR-MDP, with $g^{\star}_{\pi_{-i}}$ equal to the optimal robust average reward $g^{\star}_{\pi_{-i}}\;=\;\sup_{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\nu}_{P}.$
For each state $s\in\mathcal{S}$ and each joint policy $\nu\in\Delta(\mathcal{A}_{i})$ of player $i$, define the one-step Bellman functional as
$F^{\pi_{-i}}_{s}(\nu)\;:=\;\sum_{a_{i}\in\mathcal{A}_{i}}\nu(a_{i})\Big(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}_{\pi_{-i}}+\sigma_{(\mathcal{P}^{\pi_{-i}})_{s,a_{i}}}(u^{\star}_{\pi_{-i}})\Big),$ (9)
where $r^{\pi_{-i}}_{i}(s,a_{i})=\sum_{a_{-i}}\pi_{-i}(a_{-i}\mid s)\,r_{i}(s,(a_{i},a_{-i}))$ and $(\mathcal{P}^{\pi_{-i}})_{s,a_{i}}$ is the induced uncertainty slice at $(s,a_{i})$. Intuitively, $F^{\pi_{-i}}_{s}(\nu)$ is the robust one-step Bellman return at state $s$ when player $i$ plays $\nu(\cdot\mid s)$ and the others follow $\pi_{-i}$. We then define the Bellman-greedy response set of player $i$ w.r.t. $\pi_{-i}$ by
$\mathcal{G}_{i}(\pi_{-i})~\triangleq~\bigl\{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i}):\;F^{\pi_{-i}}_{s}(\nu(\cdot\mid s))=\max_{\mu\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\mu),\ \forall s\in\mathcal{S}\bigr\}.$ (10)
Thus $\mathcal{G}_{i}(\pi_{-i})$ consists of stationary policies that are pointwise maximizers of the optimal Bellman right-hand side for the induced DR-MDP $M_{i}(\pi_{-i})$.
By Lemma 5.1 and strong duality for weakly communicating DR-MDPs (wang2025bellman, Thm. 2), the maxima in (9) are attained for every $s$, so $\mathcal{G}_{i}(\pi_{-i})$ is nonempty. Unlike in the irreducible case, however, the presence of transient states means that a policy greedy w.r.t. a Bellman solution need not be globally optimal; in particular, not every robust best response arises as a maximizer of (9). Nonetheless, we will show that
$\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i}),$
so $\mathcal{G}_{i}(\pi_{-i})$ is a subset of the best-response set that is still rich enough to establish existence of Nash equilibria.
Finally, define the joint map
$\mathcal{G}(\pi)\;\triangleq\;\times_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$
on the compact convex policy space $\mathcal{X}=\times_{i\in\mathcal{N}}\times_{s\in\mathcal{S}}\Delta(\mathcal{A}_{i})$. If each $\mathcal{G}_{i}$ satisfies the conditions of Kakutani’s fixed-point theorem (Lemma 4.2), then $G$ admits a fixed point $\pi^{\star}\in G(\pi^{\star})$, and the inclusion $\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i})$ guarantees that this $\pi^{\star}$ is also a fixed point of the best-response map BR and hence a robust NE.
We summarize the structural properties of the maps $\mathcal{G}_{i}$ in the next theorem; the proof is deferred to Appendix D.
Theorem 5.1 (NE of weakly communicating DR-MGs).
Suppose the average-reward DR-MG satisfies Assumption 2. Then, for every player $i\in\mathcal{N}$, the following hold:
  1. 1.
    For every stationary joint policy $\pi_{-i}$ of the other players, the set $\mathcal{G}_{i}(\pi_{-i})$ is nonempty, compact, and convex.
  2. 2.
    The set-valued map $\mathcal{G}_{i}$ is upper hemi-continuous in the product topology.
  3. 3.
    For every stationary joint policy $\pi_{-i}$, any Bellman-greedy response is a robust best response: $\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i}).$
Consequently, the joint map $\mathcal{G}(\pi)=\times_{i}\mathcal{G}_{i}(\pi_{-i})$ satisfies Kakutani’s conditions (Lemma 4.2), admits a fixed point $\pi^{\star}\in G(\pi^{\star})$, and any such $\pi^{\star}$ is a stationary robust Nash equilibrium of the DR-MG.
Our results hence imply that DR-MGs with average reward is also solvable, i.e., have stationary robust NE, even under the weakly communicating settings. Due to the complicated structures, our proof techniques are fundamentally different from the irreducible case, and are deferred to Appendix D.
@@ 6 Algorithmic Solutions
In this section, we develop algorithmic solutions for irreducible average-reward DR-MGs under ˜1. Notably, finding stationary NE is significantly challenging in general-sum games, even under non-robust settings (hu2003nash; li2007reinforcement; li2003learning; zinkevich2005cyclic).
6.1 Nash Iteration
Inspired by the widely used value iteration algorithm, we first develop our robust Nash-Iteration in Algorithm 1.
@@ 6.1 Nash Iteration
Initialization: $h_{i}=0,\forall i\in\mathcal{N}$
@@
Output: $\pi$
Value-iteration has been extensively studied in algorithmic Markov game studies (hu2003nash; li2007reinforcement; li2003learning; liu2021sharp; shi2024sample; farhat2025online). Our algorithm maintains an estimation, $Q_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). We thus adopt these assumptions and derive the convergence guarantee as follows.
@@
Under ˜1 and some additional assumptions (see Assumption 3 in Appendix), Algorithm˜1 converges to a robust NE.
6.2 Proximal Descent Algorithm
Despite the convergence guarantee we obtained in Theorem˜6.1, the algorithm suffers from two disadvantages. Firstly, the convergence is only guaranteed with additional assumptions, which are standard but may not hold generally (filar2012competitive); Moreover, in each step, Algorithm˜1 solves for a NE of a matrix game, which can be PPAD-complete in the worst case (daskalakis2009complexity; chen2009settling). Although assuming such an oracle is also standard in Markov game studies (liu2021sharp; hu2003nash; shi2024sample), yet we are motivated to find other algorithms with better computational complexity.
In this section, we propose another algorithm to address the two disadvantages. Our algorithm is based on an equivalent characterization of the robust NE, in terms of the temporal difference (TD) errors. Specifically, given a joint policy $\pi$ and agent $k$’s robust average reward and robust bias $(g_{k}^{\pi},V_{k}^{\pi})$ (unique up to constant vector, per Theorem˜4.1), define a TD error as
$\Omega_{R}^{k}(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi})\triangleq\ g_{k}^{\pi}+V_{k}^{\pi}(s)-\sum_{a_{-k}}\pi_{-k}(a_{-k}\mid s)(r_{k}(s,a_{-k},a_{k})+\sigma_{\mathcal{P}_{s}^{a_{-k},a_{k}}}(V_{k}^{\pi})).$
Note that from (6), it holds that $\sum_{a_{k}}\pi_{k}(a_{k}|s)\Omega_{R}^{k}(s,a_{k};\pi_{-k}|V_{k}^{\pi},g_{k}^{\pi})=0,\forall k,s.$ Moreover, we define
$G_{k,s}(\pi)\triangleq-\min_{a_{k}}\Omega_{R}^{k}(s,a_{k};\cdot),\quad G(\pi)\triangleq\sum_{k=1}^{K}\sum_{s\in S}G_{k,s}(\pi).$
We then develop a characterization of robust NE as follows.
Lemma 6.1.
For any policy $\pi$, $G(\pi)\geq 0$. Moreover, $G(\pi)=0$ if and only if $\pi$ is a robust NE.
This result directly implies that finding a robust NE is equivalent to solve the equation $G(\pi)=0$. Moreover, since $G(\pi)\geq 0$, and $G$ is entry-wise convex, identifying a robust NE can be reduced to a minimization problem: $\min_{\pi}G(\pi)$. Such a characterization and transform is inspired by TD-based reinforcement learning approaches (sutton2018reinforcement) and their applications in standard Markov games (sahabandu2024rl).
However, solving such a multi-agent optimization problem may still be challenging, especially due to the inherent instability of multi-agent algorithm (albrecht2024multi). To address this issue and enable stable convergence, we design a block-wise proximal descent algorithm. Specifically, we index blocks by player-state pairs $(k,s)\in\mathcal{B}\triangleq\mathcal{N}\times\mathcal{S}$, where each block variable is the probability vector $\pi_{k}(\cdot\mid s)\in\Delta(\mathcal{A}_{k})$. We fix a total order $\prec$ on $\mathcal{B}$ (e.g., lexicographic by states then players). For a block $b\equiv(k,s)$, we define the sets of earlier and later blocks:
$\mathcal{B}_{<k,s>}\triangleq\{(i,t)\in\mathcal{B}:\ (i,t)\prec(k,s)\},$ (11)
$\mathcal{B}_{>k,s>}\triangleq\{(i,t)\in\mathcal{B}:\ (i,t)\succ(k,s)\}.$ (12)
Our algorithm is then updated block-wisely. Namely, when update for a block, we will set the earlier block variables to the updated ones, and apply a one-step proximal descent (i.e., mirror descent) to the optimistic estimation of $G$, $\widehat{G}_{n}\big(\pi^{n+1}_{<k,s>},\mu,\pi^{n}_{>k,s>}\big)$ (see its definition in (114)). We note that such a design is similar to the optimistic iteration studied in two-player zero-sum games (wei2021last; cen2022faster), and is shown to be more stable than simultaneously updating (daskalakis2020independent; nekoei2023dealing). We then present our algorithm in Algorithm˜2. Our algorithm is inspired by the actor-critic structure in reinforcement learning (sutton2018reinforcement).
Algorithm 2 Proximal Descent for DR-MGs
1:Initialization: $\pi=\pi^{0}$, stepsizes $\{\eta_{n}\}_{n\geq 0}$.
2:for $n=0,1,2,\dots$ do
3:  For each $k$, solve (6) at $\pi^{n}$ to obtain $(g_{k}^{n},V_{k}^{n})$.
4:  for $(k,s)\in\mathcal{B}$ do
5:   $\pi^{n+1}_{k}(s)\ \in\ \arg\min_{\mu\in\Delta(\mathcal{A}_{k})}\big\{\frac{1}{2\eta_{n}}\,\big\|\mu-\pi^{n}_{k}(s)\big\|_{2}^{2}+\widehat{G}_{n}\big(\pi^{n+1}_{<k,s>},\mu,\pi^{n}_{>k,s>}\big)\big\}$
6:  end for
7:end for
Notably, each step of our Algorithm˜2 can be executed in a polynomial complexity. Specifically, solving (6) can be done with polynomial complexity, e.g., (xu2025finite); The proximal update step is a convex optimization per block, which is also polynomial-time. Compared to Algorithm˜1 which requires a NP-hard subroutine, our algorithm is more practical and feasible.
We then derive our convergence guarantee merely under ˜1, as follows.
Theorem 6.2.
Set $\sum_{n}^{\infty}\eta_{n}<\infty$. Then $G(\pi^{n})$ converges: $\lim_{n\to\infty}G(\pi^{n})=0,$ and $\pi^{n}$ weakly converges to a robust NE, i.e., every cluster point $\pi^{\star}$ of $\{\pi^{n}\}$ is a robust NE. Moreover, if the DR-MG only has finitely many robust NE, then $\{\pi^{n}\}$ converges to some NE.
We show that, with a suitable choice of stepsizes, the TD error converges to 0. We note that this convergence does not imply convergence of $\pi^{n}$, as $\pi^{n}$ can be cycling among distinct equilibria, while $G(\pi^{n})\to 0$. However, we show that any convergent subsequence of $\{\pi^{n}\}$ converges to a robust NE, and constructing such a subsequence is straightforward in fixed-point theory and monotone operator theory, by adopting techniques like Halpern Iteration (halpern_fixed_1967) or Tikhonov Regularization.
Remark 3.
Although our Algorithm˜2 enjoys a slightly weaker convergence guarantee than Algorithm˜1, it does not require any additional assumptions on the game structure or NE-computing oracle, and can be executed in a polynomial complexity. Our algorithm hence stands for the first practical algorithm for average reward DR-MGs with provably convergence guarantees.